OpenStudy (anonymous):

If f(x)=x+5 and g(x) =x^2-3, find the following, g(f(x))= f(f(-5))= f(f(x))=

OpenStudy (stamp):

\[f(x)=x+5,\ g(x)=x^2-3\]\[f(x^2-3)=(x^2-3)+5=x^2+2\]Next\[f(-5)=(-5)+5=0\]\[f(0)=0+5=5\]Lastly\[f(x+5)=(x+5)+5=x+10\]

OpenStudy (stamp):

\[g(f(x))=(x+5)^2-3\]

OpenStudy (anonymous):

I see how the its done, thanks.